Home
Present: Jack, Jeff M, Tony, John W, Rick, Dick S
Contents
1. Agenda Item III 7 IPS will be updated see attachment to these minutes After the meeting Stefan wrote I have updated the IPS to reflect the latest information from Antonio Assembly and Testing of 2 Holography transmitters was split into two separate tasks The dates should reflect our discussion with one exception Vacation time of Front End staff was not accounted for in the delivery schedule and that puts the shipment out 1 more week This may slip the schedule described above by 1 week Ongoing See Agenda Item III 8 Although the baseline is for holography at the ATF to use the AEC antenna this is a somewhat fluid situation which Dick will watch very closely keeping all informed Dick will make the definitive pronouncement as to which antenna will be used on the rx shipping date currently scheduled for August 28 9 Darrel will organize the next meeting currently set for Thursday August 10 at 15 30 UTC If you ve read this far you ll know it happened Next meeting will be in 2 week on August 24 at 15 30 UTC AGENDA ITEM II Holography Specifications and Evaluation Criteria Are the following correct The specs are from Table 2 of an Overview by Larry D Addario for an October 2004 review of ALMA holography However the original requirement for variable frequency has been removed See http www tuc nrao edu demerson osfholo larry_overview pdf Measurement error lt 10 microns rms goal lt 5 micron
2. move the system to Chile in a crunch then if at all possible allow CIPT to do the integrated user test before shipping to Chile Then CIPT can work on bug fixes during shipment and install the final version of the software in Chile just before AIV acceptance Debra requested CIPT would prefer to have the optical pointing tests and delivery to AIV be delayed e g not at the same time as holography Debra suggested February Dec or Jan would be another possibility Dick noted after the meeting that the current PSI schedule is to install the production FEs in Feb so the ATF would be mostly free for OPT testing especially if CIPT works mostly at night Nick noted that he may not be able to deliver the OPT rack in Oct Jeff Mangum noted that this has to get done before Feb because it will be needed first for Vertex checkout in Chile Nick will provide an updated schedule in a few weeks for the OPT rack delivery to the ATF Wait to see what this is before we decide on the fate of the optical pointing delivery to AIV ACTION ITEMS 1 Antonio will deliver a user manual and ICD with monitor control points to be shipped by the time the receiver is shipped currently scheduled for August 28 2 Antonio will organize a small working group to discuss holography feed issues 3 Dick will communicate with the Peter and the site IPT over the tower specs Peter will follow the ATF tower base shelter implementation with a view to dupl
3. the OSF Brian agreed that a delivery of Dec for the complete ICD would be ok Jeff Kern asked for receiver documentation Antonio said he could regenerate notes on the holography receiver which would in effect be an interim receiver manual The algorithms for receiver tuning should be available from Christian Martin should also be consulted This documentation would be delivered with Receiver 1 The holography system is currently designed for 230 V AC Not everything is dual voltage It was agreed by Dick to provide a separate 230 V transformer at the ATF where necessary and not to modify the holography system at all There was discussion about Labview and its place at the OSF For holography at the OSF all control functions will be taken care of by the CIPT receiver tuning will be in a second stage after the documentation is delivered as will archiving of monitor data However Labview will still be needed for real time monitoring of the system which although not essential will probably be convenient Brian and Dick agreed between themselves and Pedro Ramirez they would be responsible for ensuring that a PC with compatible installed Labview would be available at the OSF for this purpose Jeff Mangum pointed out that there had been a problem running Labview and CIPT software simultaneously conflicts arrive if both try to access the same monitor point simultaneously Brian agreed this would be resolved if it is still a problem with the cu
4. transmitter pointing Verify receiver complex output and stability Margin e2e tests of the system Verfy mount movement and synch with data Verify writing reading from archive Perform 1st raster and create database Confirm data analysis and display are working correctly Produce first holo image measure adjust focus and illumination Margin Science Validation a few more images Move antenna panel to calibrate Duration 322 days 30 wks 32 wks 30 wks 10 wks 10 days 5 days 53 days 40 days 4 days 1 day 3 days 0 days 21 days 1 day 7 days 3 days 5 days 5 days 5 days 3 days 0 days 5 days 2 days 8 days 1 day 1 day 1 day 1 day 4 days 29 days 5 days 1 day 2 days 2 days 5 days 2 days 12 days 18 days 6 days 2 days Finish Aug 06 Tue 01 11 05 Thu 25 01 07 Tue 01 11 05 Mon 29 05 06 Tue 01 11 05 Mon 12 06 06 Tue 01 11 05 Mon 29 05 06 Mon 13 03 06 Mon 13 03 06 Mon 07 08 06 Fri 19 05 06 Fri 24 03 06 Fri 11 08 06 Mon 19 06 06 Wed 30 08 06 Mon 19 06 06 Fri 11 08 06 Mon 21 08 06 Thu 24 08 06 Fri 25 08 06 Fri 25 08 06 Mon 28 08 06 Wed 30 08 06 Wed 30 08 06 Wed 30 08 06 Mon 14 08 06 Mon 14 08 06 Mon 21 08 06 Thu 31 08 06 Tue 12 09 06 Mon 14 08 06 Tue 29 08 06 Tue 05 09 06 Sep 06 Oct 06 30 08 Mon 14 08 06 Fri 18 08 06 Mon 14 08 06 Thu 31 08 06 Thu 07 09 06 Fri 08 09 06 Tue 12 09 06 i Mon 28 08
5. 06 Mon 28 08 06 s Mon 28 08 06 Fri 01 09 06 TE Fri 08 09 06 Mon 11 09 06 Wed 13 09 06 Wed 13 09 06 Thu 14 09 06 Fri 15 09 06 Mon 18 09 06 Tue 19 09 06 Fri 22 09 06 Wed 13 09 06 Thu 14 09 06 Fri 15 09 06 Mon 18 09 06 Fri 22 09 06 Mon 25 09 06 Thu 02 11 06 Mon 25 09 06 Fri 29 09 06 Mon 02 10 06 Mon 02 10 06 Tue 03 10 06 Wed 04 10 06 Thu 05 10 06 Fri 06 10 06 Mon 09 10 06 Fri 13 10 06 Mon 16 10 06 Tue 17 10 06 Wed 18 10 06 Thu 02 11 06 Fri 03 11 06 Tue 28 11 06 Fri 03 11 06 Mon 13 11 06 Fri 10 11 06 Tue 14 11 06 Fri 18 08 06 LW Nov 06 30 _ 06 13 20 27 03 10 17 24 04 08 15 22 29 05 12 19 26 03 10 17 Dec 06 Project holgraphy at ATF Date Fri 11 08 06 Task Split Milestone Progress Project Summary Vy Summary External Tasks External Milestone Deadline Page 1 ID Task Name Duration Finish Aug 06 Sep 06 Oct 06 Nov 06 Dec 06 6 30 06 13 20 27 03 10 17 24 01 08 15 22 29 05 12 19 26 03 10 17 Integrated User Tests Wed 15 11 06 Tue 21 11 06 i L 48 Test Readiness Review 1day Wed 22 11 06 Wed 22 11 06 49 Software rework for high priority problems 5days Wed 22 11 06 Tue 28 11 06 50 51 Acceptance tests
6. 2days Wed 29 11 06 Thu 30 11 06 52 Margin 5 days Fri 01 12 06 Thu 07 12 06 53 Pack up 5 days Fri 08 12 06 Thu 14 12 06 h 54 ship to Chile 20 days Fri 15 12 06 Thu 11 01 07 55 EJ Reconfigure sofware install test functionality in lab 10 days Fri 12 01 07 Thu 25 01 07 56 Holography system ready for installation in first antenna 0days Thu 25 01 07 Thu 25 01 07 Project holgraphy at ATF Task COS Progress Summary gy External Tasks Deadline Jb Page 2
7. Minutes of Teleconference on ALMA ATF amp OSF Holography Planning Thursday August 10 2006 15 30 UTC Minutes by DTE last changed 2006 08 15 Participants Beasley part of meeting Brito Emerson Glendenning Kern Lucas Mangum Murowinski Ocampo Perfetto Ramirez Shepherd Sramek All future meetings will use the same call in details From USA 866 814 1347 Outside USA 1 517 444 3243 Participant Passcode 3155752 Leader Passcode 1874599 Minutes of our last 2006 07 20 meeting are at http www tuc nrao edu demerson osfholo mins2006 07 20 1 1 attach pdf AGENDA 2006 08 10 I Old Action Items Il Holography Specifications and Evaluation Criteria HI Schedule AGENDA ITEM I Old Action Items See minutes of last meeting at http www tuc nrao edu demerson osfholo mins2006 07 20 1 1 attach pdf 1 Antonio to give Brian revised ICD It is understood that the exact ranges may change as the design progresses however the M C points and formats can be fixed This is not yet completed however it is not critical Brian would have preferred to have received the complete information but given that the ICD only adds M C points i e it is backwards compatible it can wait for 2 months Jeff Kern agreed Receiver 2 will be delivered late December 2006 with holography transmitter 2 Everyone agreed these should be sent to the ATF as spares where they may be used to help possible and eventual debugging at
8. dule For Darrel s item III I put together a picture of a schedule of work at ATF on the holography tests to use as a strawman for the discussion tomorrow This plan is extracted from a recent discussion between a few of you who are trying to plan that work out PDF and MPP versions of the same information are attached here There was considerable discussion and Rick agreed to produce a revised schedule This is attached to the current minutes below One issue was clarification needed as to which cables and cable work corresponded to which line item There was discussion as to whether we should be planning for a 5 day or 7 day working week and the relationship between this and contingency time in the schedule The following are extracted from notes written by Debra CIPT needs the following items added to the holography schedule Duration Task 1 week Integrated user tests at the ATF 1 week Bug fixes and verification at the ATF for the highest priority bugs that must be fixed before acceptance by AIV CIPT would prefer to have this scheduled after the Science validation tasks Rick asked What if we deliver the system to Chile without these tests and then do the tests in Chile AIV acceptance rules Debra would have to be bent for this CIPT would stongly prefer not to have this happen since AIV folks will have to deal with bugs in the system at first and not a complete and documented system Not good If you have to
9. icating it at the OSF 4 Dick Brian and Pedro will ensure the right Labview and PC get delivered to the OSF 5 Dick will pronounce on which antenna on August 28 or on whatever the rx shipment date is if that slips 6 Rick will produce a revised schedule and distribute see attachment below 7 Darrel will organize the next teleconference for Thursday August 24 at 15 30 UTC 5 TPRP 9 l N o Aal A oj N 11 La Li 12 13 14 15 16 17 La Li 18 19 20 EN 21 22 23 24 mai a a a ma a ar a ma ar E a E a a ao a a a o gt 44 45 46 Task Name Holography System Develop transmitter electronics Repackage holography receiver Develop ancillary electronics cabling power supply ABM TE and 25MHz etc Develop control software Modify reduction software for data format Ship Tx cables and spare minirack to AOC NTC Rx and Tx Integrate test tx rx at CV System test at NTC Pack Ship to AOC Receive Tx and Rx at AOC AOC and ATF preparation Inspect Tower Install Tx cables Install Tx at ATF and verify control Prepare RF cables Ensure 230V power is available at ATF Control software checkout at AOC Upgrade software at ATF Antenna VA AEC decision prepare antenna for electronics Install receiver Checkout Check M C and 1st point of Tx Ensure Tx and Rx on correct frequency confirm adequate received power Peak
10. rrent software 2 Antonio will organize a small working group to consider the holography feed issues what feeds are or are not acceptable and what timescale for delivery of feeds to ATF and or OSF This working group has not yet been formed but will be soon There was discussion as to which feeds were in Robert Ridgeway s hands Further discussion of feed acceptability would take place within this new working group 3 Dick will keep contact with site over the tower specs to ensure that adequate equipment haulage and a platform are included This is not thought to be a big issue but Dick will keep contact with Peter and SiteIPT over the tower specs 4 Site IPT should discuss with Jack the needs for the shelter at the OSF tower base Probably nothing more to do here Peter will see what is put in place at the ATF during testing there and duplicating something quickly for the OSF later 5 Rodrigo will put details of the holography rack onto EDM See attachment to these minutes Rodrigo comments These see attachment are a few general assembly drawings I have for documentation They can be used as a reference In the binder that will be delivered together with the mini rack there will also be detailed PCB drawings data sheets test results the SoW purchase orders and the disc configuration for the Ethernet Power Strip Completed 6 Dick and Brian will circulate an agreed schedule for the ATF activities Ongoing See
11. s rms Transverse resolution 0 1 m or better Measurement time 60 mins goal 30 mins Frequency 80 120 GHz fixed frequency What component of the error is allocated to the dish feed How do we evaluate Compare the September 2006 holography measurement to 2004 measurement How close should they be What are the pass fail criteria Jeff Mangum commented that in practice a transverse resolution of 0 2m rather than 0 1 has proved adequate The 10 micron rms and 5 micron goal are repeatability errors not including systematic effects such as main dish feed phase pattern As a holography verification criterion it was agreed that new holography maps should be compared to the previous 2004 holography data A correction should be applied for temperature differences but after that correction the rms of the difference map between new 2006 data and 2004 data should be less than 14 microns If the holography test failed this 14 micron criterion then we need to know if the problem is with holography measurement or with the antenna changing shape In such circumstances the holography system might be moved to the other antenna and fresh measurements made Evaluation of those results could determine if holography is the problem or much more seriously if an antenna has changed shape with time AGENDA ITEM III Schedule What are the latest dates Rick distributed before the meeting the following email dated 2006 08 09 with an attached sche
Download Pdf Manuals
Related Search
Related Contents
Ankle Wrap Application MP3X17 Manual - Chassis Plans Caractérisation de couches épaisses de semi Copyright © All rights reserved.
Failed to retrieve file